OpenStudy (lilai3):

Please help! Will award medal! A fitness center has set a goal to have 500 members. The fitness center already has 150 members and adds an average of 25 members per month,. The function f(x)=25x+150 represents the membership after x months. Graph the function to determine the number of months it will take for the fitness center to reach its membership goal.

OpenStudy (stamp):

think like a programmer, with algebra as your language define your terms goal = 500 membership = 150 + 25*month Find month for membership to equal goal. membership = goal 150 + 25*month = 500

OpenStudy (lilai3):

so uhm... thanks?

OpenStudy (stamp):

150 + 25*month = 500 25*month = 350 month = 350/25 month = 14